in one half year, the earth moves halfway around its orbit, a circle of radius 1.50×10^11m centered on the sun . what is the average speed and what is the magnitude of the average velocity for this time interval


1
Expert's answer
2021-02-17T11:01:21-0500

1)


"v=\frac{\pi r}{t}\\v=\frac{\pi (1.5\cdot10^{11})}{0.5(365)(24)(3600)}=3.0\cdot10^{4}\frac{m}{s}"

2)


"v_{av}=0.5(v+(-v))=0\frac{m}{s}"